Solve (v+1)^2-72=0 , where v is a real number.
Simplify your answer as much as possible.
If there is more than one solution, separate them with commas.

Answer by Alan3354(69443)

(v+1)^2-72=0
(v+1)^2 = 72
v+1 = ± sqrt(72)
v = -1 ± 6sqrt(2)